Heidelberg weekly meeting

Europe/Zurich
03.410 (Physikalisches Institut, Heidelberg)

03.410

Physikalisches Institut, Heidelberg

Description

We will have the meeting via Zoom: https://cern.zoom.us/j/92591061032?pwd=MkZzY2g1cG1YZGFwK244c210dk44dz09

Note: this is a temporary room until we get the institute account.

We will keep HeiCONF as a backup: https://heiconf.uni-heidelberg.de/sch-nqu-dd2   Access Code: 858824

Vidyo will be used as a back-up. Room extension 10823320 or use this link.

 

Tentative minute taker list at https://docs.google.com/spreadsheets/d/15oo_a_1WPxPBOtlH0LDEKG1128y1O2lfswpUAvPCQOY/edit#gid=0

    • 2:00 PM 2:45 PM
      Round Table 45m

      Reports of work progress

      Speakers: Adrian Herkert (Physikalisches Institut, Uni Heidelberg) , Adrian Stoll, Alena Weber (Universität Heidelberg) , Andre Schoening (University Heidelberg, Institute of Physics) , Annie Meneses Gonzalez (PhD Student) , Benjamin Weinläder (PI Heidelberg) , David Maximilian Immig (Universität Heidelberg) , Dohun Kim (Physikalisches Institut University of Heidelberg) , Florian Frauen, Frank Meier Aeschbacher (PSI - Paul Scherrer Institut) , Heiko Augustin (PI Uni Heidelberg) , Jakob Stricker, Lars Noehte (PI Heidelberg) , Luigi Vigani (University of Heidelberg) , Lukas Mandok (Physikalisches Institut Heidelberg) , Marius Menzel (Institute of Physics, University of Heidelberg) , Sebastian Dittmeier (Physikalisches Institut, Universität Heidelberg) , Sebastian Preuß (PI Uni Heidelberg) , Stephan Lachnit, Thomas Theodor Rudzki (Physikalisches Institut Heidelberg)

      Annie: at testbeam. Luigi: confirmed 90% efficiency for 50-um-thin 20-ohm-cm Mupix10? Yes. But maybe not enough statistics for in-pixel efficiency, gotta check. With the same setup we obtain ~99.5% efficiency with 200-ohm-cm Mupix10, so setup should be fine. Heiko: with Mupix7 tuning did not improve efficiency that much, but got a larger plateau. Note: the n-well design should be optimized for large resistivity (small n-well size). It is possible that 20 ohm*cm are not depleted at the corners. Lennhart and Adrian also checked the sensor-to-sensor variations. Worth cross-check. Got designs from Ivan to check.


      Dohun: testbeam. Telescope installed. DUT AtlasPix3.0. AtlasPix3.1 did not work properly in testbeam area (high 8/10bit error rate). Scanning threshold at different thresholds. Going to align telescope afterwards then measure. Going to check only irradiated AtlasPix3.

      Heiko: check: the distance from n-well to pixel corner went up from Mupix7 to Mupix8 and Mupix10. 20 Ohm cm Mupix10 migh hot have lateral depletion then. Expected inefficiency at the corners. Suggestion to Annie: start analysis with 200 Ohm cm low voltage cause we know it is from edge effects.
      Mupix10 from Optim picked last week. Gotta be checked with probe station next week, hoping plasma etching works. By eye no big difference observed from previous chips (still could see directional polishing). If no improvements re-try with atlaspix3.1 (65 V breakdown) and DISCO (same as ALICE). Also, ALICE TowerJazz technology is different, so no guarantee the same process works. We might need more fancy processing (p-dope implant on back...?)

      Thomas: started writing thesis. Still investigating difference in temperature readings. Asked Mainz for checking in a climate chamber. Heidelberg can do. Luigi will check.
      Trying to bond single Mupix10 chip to single flex. Really difficult to bond to insert. HDIs are rigid. If new inserts needed coming to HD anyway on Monday. We have to decide what to bring back from PSI, otherwise we should pay an extra for importing materials. Also, to bring back all stuff on balcony check with MEG.

      Jan: delays. Bonder was sick. AtlasPix3.0 removed from gel-pack. AtlasPix3 sensors were forgot by Dohun, gonna be done by Thomas on Monday. Mupix10 inserts produced. Spare waffle packs would be nice. Heiko: I will prepare the box with AtlasPix3 chips. Dohun should also have a CPU cooling block connected to Peltier for cooling the air, it should be sent as well. Run2020 insert design in bitbucket.

      Luigi: cleaned the desk. Working on paper with Cristina. Gonna check with Ralf probe station availability on Monday. Debugging probe-card now the main challenge these weeks.

      Thomas: how should I improve bonding to insert pads? Heiko: try and increase pressure until you see bending on gold pads. Will keep checking.
       

    • 2:45 PM 2:55 PM
      AOB